OpenStudy (anonymous):

1/36 - 1/x^2 over 1/6 + 1/x a. 6-x/6x b. 6x/x-6 c. x-6/6x d. 6x/6-x

OpenStudy (anonymous):

@ankit042

OpenStudy (luigi0210):

Get common denominators

OpenStudy (ranga):

I will simplify the numerator. You can follow the same procedure for the denominator and then simplify. 1/36 - 1/x^2 = (x^2 - 36) / (36x^2) = (x+6)(x-6) / (36x^2) Denominator = 1/6 + 1/x = ?

OpenStudy (anonymous):

@ankit042